OpenStudy (minatsukisaya95):

how to find the solution in the interval of 0 to 360 deg for 2cos u+cos 2u=0

OpenStudy (lncognlto):

\(\Huge\bf \color{black}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{black}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{#ff7100}{Welcome~to~\color{#00B4ff}{\underline{\overline{Open}}}\color{#7cc517}{\underline{\overline{Study!!}}}}\)

OpenStudy (debbieg):

I think if you make a substitution using the identity: \(\Large \cos2u=2\cos^2u-1\) Then you will have an equation of quadratic type and can solve with quadratic equation methods.

OpenStudy (minatsukisaya95):

i have done that but it is not real number

OpenStudy (debbieg):

You should get 2 solutions for cos(u).... then use the inverse function to find u. Can you show your work to solve the quadratic? You should get 2 irrational solutions.

OpenStudy (minatsukisaya95):

yep i get two irrational solution from that how do i do next

OpenStudy (debbieg):

\[\Large 2\cos u+\cos 2u=0\]\[\Large 2\cos u+2\cos^2u-1=0\] disriminant =\(\large b^2-4ac=4-4(2)(-1)=4+8=12\) so should have 2 irrational solutions.

OpenStudy (debbieg):

OK, remember that those are solutions for cosu. So you have cos(u)=a You want u. and you know the cosine of u. What "tool" (think of a function that you can use) takes the COSINE VALUE of an angle as the input, and gives you the ANGLE MEASURE as the output?

OpenStudy (debbieg):

These are your solutions for the quadratic? \[\Large \cos u=\frac{ -1\pm \sqrt{3} }{ 2}\]

OpenStudy (minatsukisaya95):

oh ok so the two solution is the exact value of cos u ok ok i got it thank u

OpenStudy (debbieg):

you're welcome. :)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."